# A Survey on Integration of Large Language Models with Intelligent Robots

### v2

@@ 7.3 Planning Prompt: Few-shot Planning
The task descriptions include task objectives, expected behaviors, and response formats similar to conventional prompts. However, unlike previous ones, this prompt specifies the robot’s constraints, including initial states and action limitations-constraints previously unaddressed. For example, the term “CANNOT” in Table 3 emphasizes the robot’s limitation to manipulating only one object per action. Moreover, these constraints extend to the rules governing the “Done” action, indicating task completion.
The examples demonstrate input-output pairs that guide the LLM in generating the desired action. The examples adapt the generic “object” argument in the allowed actions (e.g., “Close (object)”) to specific object names such as “drawer” or “paper.” reinforcing task constraints written in the task description. For instance, the second example returns the “Done” signal instead of further planning after achieving the task objective.
The target context provides the current scenario, including the task, allowed actions, visible objects, executed plans, and a next plan, same as examples. We allow the LLM to fill in the blank space after “next plan:” suggesting the next action without adding unnecessary elements like line breaks, ensuring output precision.
@@ 7.4 Code-Generation Prompt: Reward Design
Task descriptions define the expected robot behavior and task conditions for the LLM, including the robot’s control strategies and the action space of the two-joint robot arm. We particularly specify the action space as a continuous “Box” space using an API from Gymnasium, assuming the LLM’s familiarity with well-known library functions. Then, this description leads the LLM to grasp the overarching RL objective of the defined actions.
@@
Goals and constraints provide the task objectives and limitations that guide the reward contents. We clearly define the initial setup, goal assignment, and goal conditions, aiming to exclude unnecessary reward components, such as penalizing high velocities for smooth motion. Note that we recommend the use of concise and consistent words, such as “torque”, as used in the task descriptions, instead of “power,” though the linguistic meaning is similar.
@@ 8 Conclusion
In this survey, we have investigated the current robotics research works with large language models in terms of intelligent robot components encompassing communication, perception, planning, and control. This component-wise investigation reveals how researchers integrate LLMs to overcome challenges inherent in pre-LLM approaches across various tasks, thereby offering a comprehensive understanding of LLMs’ impact in this field. Within each component area, we examine the improvement of methodologies proposed to maximize the utilization of LLMs’ capabilities and enhance the integrity of their responses. Additionally, our survey offers guidelines for prompt engineering in each component area, supplemented with key examples of prompt components, to provide practical insights for researcher entering this field. The core contribution of this paper is to highlight the transformative impact of LLMs in robotics, enabling the development of versatile and intelligent robots with limited resources.

### v3

@@ 7.3 Planning Prompt: Few-shot Planning
The task descriptions include task objectives, expected behaviors, and response formats similar to conventional prompts. However, unlike previous ones, this prompt specifies the robot’s constraints, including initial states and action limitations. For example, the term ‘CANNOT’ in Table 3 emphasizes the robot’s limitation to manipulating only one object per action. Moreover, these constraints extend to the rules governing the ‘Done’ action, indicating task completion.
The examples demonstrate input-output pairs that guide the LLM in generating the desired action. The examples adapt the generic ‘object’ argument in the allowed actions (e.g., ‘Close (object)’) to specific object names such as ‘drawer’ or ‘paper,’ reinforcing task constraints written in the task description. For instance, the second example returns the ‘Done’ signal instead of further planning after achieving the task objective.
The target context provides information about the current scenario, including ‘task,’ ‘allowed actions,’ ‘visible objects,’ ‘executed plans,’ and ‘next plan,’ as shown in the examples. We allow the LLM to fill in the blank space after ‘next plan:’ suggesting the next action without adding unnecessary elements like line breaks, ensuring output precision.
@@ 7.4 Code-Generation Prompt: Reward Design
Task descriptions define the expected robot behavior and task conditions for the LLM, including the robot’s control strategies and the action space of the two-joint robot arm. We particularly specify the action space as a continuous ‘Box’ space using an API from Gymnasium, assuming the LLM’s familiarity with well-known library functions. Then, this description leads the LLM to grasp the overarching RL objective of the defined actions.
@@
Goals and constraints provide the task objectives and limitations that guide the reward contents. We clearly define the initial setup, goal assignment, and goal conditions, aiming to exclude unnecessary reward components, such as penalizing high velocities for smooth motion. Note that we recommend the use of concise and consistent words, such as ‘torque,’ as used in the task descriptions, instead of ‘power.’ This ensures the generated reward function aligns with the specified task requirements without introducing ambiguities or unintended penalties.
@@ 8 Conclusion
In this survey, we have investigated the current robotics research works with large language models in terms of intelligent robot components encompassing communication, perception, planning, and control. This component-wise investigation reveals how researchers integrate LLMs to overcome challenges inherent in pre-LLM approaches across various tasks, thereby offering a comprehensive understanding of LLMs’ impact in this field. Within each component area, we examine the improvement of methodologies proposed to maximize the utilization of LLMs’ capabilities and enhance the integrity of their responses. Additionally, our survey offers guidelines for prompt engineering in each component area, supplemented with key examples of prompt components, to provide practical insights for researcher entering this field. The core contribution of this paper is to highlight the transformative impact of LLMs in robotics, enabling the development of versatile and intelligent robots. By synthesizing these insights, we aim to guide future research on integrating LLMs into robotic system.
[…]
Acknowledgements This research was supported by the DRB-KAIST SketchTheFuture Research Center and the KAIST Convergence Research Institute Operation Program.